OpenStudy (anonymous):

Is the line through the points R(-1,3) and S(2,-7) parallel to the graph of the line given by the equation, 10x+3y=6? A. No, the slope have opposite signs B. Yes, the slopes have the same sign and value. C. Yes, the lines both decrease to the right. D. No, the lines have unequal slopes.

OpenStudy (raden):

check the values slope of both line, if they are same then both line is parallel the slope of line which through points R(-1,3) and S(2,-7) is : m = (y2 - y1)/(x2 - x1) = (-7 - 3)/(2 - (-1)) = -10/3. while the line of 10x+3y=6 has the slope m = -a/b = -10/3 m1 = m2 = -10/3. (yes, they are parallel)

OpenStudy (anonymous):

OH thank you @RadEn , u save me thanks

OpenStudy (raden):

wlcm :)